\begin{thebibliography}{10}

\bibitem{hervik}
{\O}.~Gr{\o}n and S.~Hervik, {\em Einstein's General Theory of Relativity: With
  Modern Applications in Cosmology}.
\newblock Springer, first~ed., 2007.

\bibitem{elliscosmo}
G.~F. Ellis, R.~Maartens, and M.~A. MacCallum, {\em Relativistic Cosmology}.
\newblock Cambridge University Press, first~ed., 2012.

\bibitem{2007Prama..69...15E}
G.~F.~R. {Ellis}, ``{On the Raychaudhuri equation},'' {\em Pramana}, vol.~69,
  p.~15, July 2007.

\bibitem{tolmanward}
R.~C. Tolman and M.~Ward, ``On the behavior of non-static models of the
  universe when the cosmological term is omitted,'' {\em Physical Review},
  vol.~39, pp.~835--843, 1932.

\bibitem{raych1955}
A.~Raychaudhuri, ``Relativistic cosmology. {{I}},'' {\em Physical Review},
  vol.~98, pp.~1123--1126, 1955.

\bibitem{1931Natur.128..704L}
A.~G. {Lema{\^i}tre}, ``{Contributions to a British Association Discussion on
  the Evolution of the Universe.},'' {\em \nat}, vol.~128, pp.~704--706, Oct.
  1931.

\bibitem{1931Natur.127..706L}
G.~{Lema{\^i}tre}, ``{The Beginning of the World from the Point of View of
  Quantum Theory.},'' {\em \nat}, vol.~127, p.~706, May 1931.

\bibitem{1970Natur.228..544H}
F.~{Hoyle} and J.~V. {Narlikar}, ``{Effect of Quantum Conditions in a Friedmann
  Cosmology},'' {\em \nat}, vol.~228, pp.~544--545, Nov. 1970.

\bibitem{1985GReGr..17..397E}
G.~F.~R. {Ellis} and W.~L. {Roque}, ``{The nature of the initial
  singularity},'' {\em General Relativity and Gravitation}, vol.~17,
  pp.~397--406, Apr. 1985.

\bibitem{1985GReGr..17..769M}
V.~{Mueller} and H.-J. {Schmidt}, ``{On Bianchi type-I vacuum solutions in R +
  R-squared theories of gravitation. I - The isotropic case},'' {\em General
  Relativity and Gravitation}, vol.~17, pp.~769--781, Aug. 1985.

\bibitem{1986LIACo..26..121S}
P.~{Spindel} and P.~{Vandepeutte}, ``{Numerical investigations in
  cosmogenesis},'' in {\em Liege International Astrophysical Colloquia},
  vol.~26 of {\em Liege International Astrophysical Colloquia}, pp.~121--131,
  1986.

\bibitem{1988PNAS...85.7428P}
I.~{Prigogine}, J.~{Geheniau}, E.~{Gunzig}, and P.~{Nardone}, ``{Thermodynamics
  of Cosmological Matter Creation},'' {\em Proceedings of the National Academy
  of Science}, vol.~85, pp.~7428--7432, Oct. 1988.

\bibitem{1991GReGr..23..527H}
R.~N. {Henriksen} and K.~{Patel}, ``{Null charts and naked singularities in
  spherically symmetric, homothetic spacetimes},'' {\em General Relativity and
  Gravitation}, vol.~23, pp.~527--581, May 1991.

\bibitem{1992ApJ...392..385R}
A.~{Rebhan}, ``{Large-scale rotational perturbations of a Friedmann universe
  with collisionless matter and primordial magnetic fields},'' {\em \apj},
  vol.~392, pp.~385--393, June 1992.

\bibitem{1993RSPSA.443..493N}
R.~P.~A.~C. {Newman}, ``{On the Structure of Conformal Singularities in
  Classical General Relativity. II Evolution Equations and a Conjecture of K.
  P. Tod},'' {\em Proceedings of the Royal Society of London Series A},
  vol.~443, pp.~493--515, Dec. 1993.

\bibitem{1971CMaPh..20...57M}
M.~A.~H. {MacCallum}, ``{A class of homogeneous cosmological models III:
  Asymptotic behaviour},'' {\em Communications in Mathematical Physics},
  vol.~20, pp.~57--84, Mar. 1971.

\bibitem{1994PhLA..188..130C}
S.~{Capozziello}, R.~{de Ritis}, and P.~{Scudellaro}, ``{Nonminimal coupling
  and cosmic no-hair theorem},'' {\em Physics Letters A}, vol.~188,
  pp.~130--136, May 1994.

\bibitem{1996Prama..47...41A}
{Abdussattar} and R.~G. {Vishwakarma}, ``{A model of the universe with decaying
  vacuum energy.},'' {\em Pramana}, vol.~47, pp.~41--55, July 1996.

\bibitem{2003PhRvD..68f3511S}
M.~{Szyd{\l}owski} and A.~{Krawiec}, ``{Dynamical system approach to
  cosmological models with a varying speed of light},'' {\em \prd}, vol.~68,
  p.~063511, Sept. 2003.

\bibitem{2004PhRvD..69l3504C}
C.~S. {Camara}, M.~R. {de Garcia Maia}, J.~C. {Carvalho}, and J.~A. {Lima},
  ``{Nonsingular FRW cosmology and nonlinear electrodynamics},'' {\em \prd},
  vol.~69, p.~123504, June 2004.

\bibitem{2005PhRvD..72f3504S}
M.~{Szyd{\l}owski}, W.~{God{\l}owski}, A.~{Krawiec}, and J.~{Golbiak}, ``{Can
  the initial singularity be detected by cosmological tests?},'' {\em \prd},
  vol.~72, p.~063504, Sept. 2005.

\bibitem{2011JETP..112...60K}
V.~V. {Karbanovski}, A.~S. {Tarasova}, A.~S. {Salimova}, G.~V. {Bilinskaya},
  and A.~N. {Sumbulov}, ``{Model of the static universe within GR},'' {\em
  Soviet Journal of Experimental and Theoretical Physics}, vol.~112,
  pp.~60--62, Jan. 2011.

\bibitem{2013IJTP...52.2991O}
G.~{Oliveira-Neto}, G.~A. {Monerat}, E.~V. {Corr{\^e}a Silva}, C.~{Neves}, and
  L.~G. {Ferreira Filho}, ``{Quantization of Friedmann-Robertson-Walker
  Spacetimes in the Presence of a Cosmological Constant and Stiff Matter},''
  {\em International Journal of Theoretical Physics}, vol.~52, pp.~2991--3006,
  Sept. 2013.

\bibitem{1994NuPhB.415..497A}
I.~{Antoniadis}, J.~{Rizos}, and K.~{Tamvakis}, ``{Singularity-free
  cosmological solutions of the superstring effective action},'' {\em Nuclear
  Physics B}, vol.~415, pp.~497--514, Mar. 1994.

\bibitem{1994CQGra..11.1919A}
J.~P. {Abreu}, P.~{Crawford}, and J.~P. {Mimoso}, ``{Exact conformal scalar
  field cosmologies},'' {\em Classical and Quantum Gravity}, vol.~11,
  pp.~1919--1939, Aug. 1994.

\bibitem{1998PhRvD..58f3504M}
G.~A. {Monerat}, H.~P. {de Oliveira}, and I.~D. {Soares}, ``{Chaos in
  preinflationary Friedmann-Robertson-Walker universes},'' {\em \prd}, vol.~58,
  p.~063504, Sept. 1998.

\bibitem{2000PhRvD..62d3526C}
A.~{Coley} and M.~{Goliath}, ``{Closed cosmologies with a perfect fluid and a
  scalar field},'' {\em \prd}, vol.~62, p.~043526, Aug. 2000.

\bibitem{2002PhRvD..66b3513I}
H.~{Ishihara}, ``{Big bang of the brane universe},'' {\em \prd}, vol.~66,
  p.~023513, July 2002.

\bibitem{2012arXiv1203.1819S}
O.-C. {Stoica}, ``{Beyond the Friedmann-Lemaitre-Robertson-Walker Big Bang
  singularity},'' {\em ArXiv e-prints}, Mar. 2012.

\bibitem{2012arXiv1207.5303C}
O.~{Cristinel Stoica}, ``{An Exploration of the Singularities in General
  Relativity},'' {\em ArXiv e-prints}, July 2012.

\bibitem{2012Entrp..14.1296G}
C.~{Gao}, ``{A Model of Nonsingular Universe},'' {\em Entropy}, vol.~14,
  pp.~1296--1305, July 2012.

\bibitem{2016arXiv160108175D}
J.~{de Haro}, J.~{Amor{\'o}s}, and S.~{Pan}, ``{A simple nonsingular
  inflationary quintessential model},'' {\em ArXiv e-prints}, Jan. 2016.

\bibitem{1986Ap&SS.123..405B}
A.~{Beesham}, ``{Comment on the Big-Bang singularity in the scale-covariant
  theory},'' {\em Astrophysics and Space Science}, vol.~123, pp.~405--407, June
  1986.

\bibitem{1982CaJPh..60..659M}
J.~W. {Moffat} and D.~{Vincent}, ``{The early universe in a generalized theory
  of gravitation},'' {\em Canadian Journal of Physics}, vol.~60, pp.~659--663,
  May 1982.

\bibitem{2007CQGra..24.6243P}
L.~{Parisi}, M.~{Bruni}, R.~{Maartens}, and K.~{Vandersloot}, ``{The Einstein
  static universe in loop quantum cosmology},'' {\em Classical and Quantum
  Gravity}, vol.~24, pp.~6243--6253, Dec. 2007.

\bibitem{2009PhRvD..79l3509K}
A.~{Kreienbuehl}, ``{Singularity avoidance and time in quantum gravity},'' {\em
  \prd}, vol.~79, p.~123509, June 2009.

\bibitem{2009PhRvD..80b3501A}
S.~{Alexander} and T.~{Biswas}, ``{Cosmological BCS mechanism and the big bang
  singularity},'' {\em \prd}, vol.~80, p.~023501, July 2009.

\bibitem{2010PhLB..688..129V}
B.~{Vakili}, ``{Classical and quantum dynamics of a perfect fluid scalar-metric
  cosmology},'' {\em Physics Letters B}, vol.~688, pp.~129--136, May 2010.

\bibitem{2012PhRvD..85b3512L}
I.~D. {Lawrie}, ``{Internal time, test clocks, and singularity resolution in
  dust-filled quantum cosmology},'' {\em \prd}, vol.~85, p.~023512, Jan. 2012.

\bibitem{2012PhRvD..85f4001P}
T.~{Paw{\l}owski} and A.~{Ashtekar}, ``{Positive cosmological constant in loop
  quantum cosmology},'' {\em \prd}, vol.~85, p.~064001, Mar. 2012.

\bibitem{2014CEJPh..12..123S}
O.-C. {Stoica}, ``{Einstein equation at singularities},'' {\em Central European
  Journal of Physics}, vol.~12, pp.~123--131, Feb. 2014.

\bibitem{2014PhRvD..90l3538P}
T.~{Paw{\l}owski}, R.~{Pierini}, and E.~{Wilson-Ewing}, ``{Loop quantum
  cosmology of a radiation-dominated flat FLRW universe},'' {\em \prd},
  vol.~90, p.~123538, Dec. 2014.

\bibitem{2016IJTP...55...71S}
O.~C. {Stoica}, ``{The Friedmann-Lema{\^i}tre-Robertson-Walker Big Bang
  Singularities are Well Behaved},'' {\em International Journal of Theoretical
  Physics}, vol.~55, pp.~71--80, Jan. 2016.

\bibitem{2007PhLB..656...96B}
M.~V. {Battisti} and G.~{Montani}, ``{The Big-Bang singularity in the framework
  of a Generalized Uncertainty Principle},'' {\em Physics Letters B}, vol.~656,
  pp.~96--101, Nov. 2007.

\bibitem{2004CQGra..21..223E}
G.~F.~R. {Ellis} and R.~{Maartens}, ``{The emergent universe: inflationary
  cosmology with no singularity},'' {\em Classical and Quantum Gravity},
  vol.~21, pp.~223--232, Jan. 2004.

\bibitem{ellis3}
S.~Hawking and G.~Ellis, {\em The large scale structure of space-time}.
\newblock Cambridge University Press, twentieth printing~ed., 2006.

\bibitem{waldbook}
R.~M. Wald, {\em General Relativity}.
\newblock University of Chicago Press, 1984~ed., 1984.

\bibitem{ellismac}
G.~Ellis and M.~MacCallum, ``A class of homogeneous cosmological models,'' {\em
  Comm. Math. Phys}, vol.~12, pp.~108--141, 1969.

\bibitem{ellis}
J.~Wainwright and G.~Ellis, {\em Dynamical Systems in Cosmology}.
\newblock Cambridge University Press, first~ed., 1997.

\bibitem{arnolddyn}
D.~Anosov, S.~K. Aranson, V.~Arnold, I.~Bronshtein, V.~Grines, and
  Y.~Il'yashenko, {\em Ordinary Differential Equations and Smooth Dynamical
  Systems}.
\newblock Springer-Verlag, third~ed., 1997.

\end{thebibliography}

\documentclass[prd, twocolumn]{revtex4}
\usepackage{amsmath}
\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}
\usepackage{wrapfig}
\usepackage{epsfig}
\usepackage{breqn}
\usepackage{amssymb}
\usepackage{subfigure}
\usepackage[usenames,dvipsnames]{pstricks}
\usepackage{epsfig}
\usepackage{pst-grad} 
\usepackage{pst-plot} 
 \usepackage[usenames,dvipsnames]{pstricks}
 \usepackage{epsfig}
 \usepackage{hyperref}
 \usepackage{pst-grad} 
 \usepackage{pst-plot} 
 \usepackage{float}

\begin{document}

\title{On Past Singularities in $k=0$ FLRW Cosmologies}
\author{Ikjyot Singh Kohli}
	\email{isk@mathstat.yorku.ca}
	\affiliation{York University - Department of Mathematics and Statistics}
\date{September 6, 2016}                                           


\begin{abstract}
The fundamental singularity theorem of FLRW cosmologies assumes that the matter content in the cosmological model obeys the strong energy condition along with a nonpositive cosmological constant which gives rise to an irrotational geodesic singularity. In this paper, we show that the important case of a spatially flat Friedmann-Lema\^{i}tre-Robertson-Walker universe with barotropic matter obeying only the \emph{weak} energy condition with a nonnegative cosmological constant also contains a past singularity. We accomplish this using topological methods from dynamical systems theory. The methods employed in this paper are sufficiently general that they could be extended to other models to study the existence of past singularities.
\end{abstract}

\maketitle

\section{Introduction}
Some cosmological solutions to the Einstein field equations are understood to begin at a certain cosmic time, which for the sake of simplicity is usually denoted as time $t = 0$, which is usually referred to as the \emph{point of time of the Big Bang} \cite{hervik}. The issue of singularities is important in the context of cosmology because this can possibly give us an answer to whether the Big Bang actually occurred or not. 

If one considers a cosmological constant denoted by $\Lambda$ and denotes the energy density and pressure of the matter in a Friedmann-Lema\^{i}tre-Robertson-Walker (FLRW) universe by $\mu$ and $p$ respectively, then the fundamental singularity theorem for FLRW cosmologies states that if $\Lambda \leq 0, \mu + 3p \geq 0$, and $\mu + p > 0$ (that is, the matter obeys the strong energy condition) in a fluid flow for which there is no acceleration or rotation, then an irrotational geodesic singularity will occur at a finite proper time  \cite{elliscosmo,2007Prama..69...15E, tolmanward, raych1955}. 

With respect to Friedmann-Lema\^{i}tre-Robertson-Walker (FLRW) spacetimes, the concept of initial singularities / big bang singularities has been studied a number of times in the literature in various contexts from classical general relativity to quantum cosmology \cite{1931Natur.128..704L, 1931Natur.127..706L, 1970Natur.228..544H, 1985GReGr..17..397E, 1985GReGr..17..769M, 1986LIACo..26..121S, 1988PNAS...85.7428P, 1991GReGr..23..527H, 1992ApJ...392..385R, 1993RSPSA.443..493N, 1971CMaPh..20...57M, 1994PhLA..188..130C, 1996Prama..47...41A, 2003PhRvD..68f3511S, 2004PhRvD..69l3504C, 2005PhRvD..72f3504S, 2011JETP..112...60K, 2013IJTP...52.2991O, 1994NuPhB.415..497A, 1994CQGra..11.1919A, 1998PhRvD..58f3504M, 2000PhRvD..62d3526C, 2002PhRvD..66b3513I, 2012arXiv1203.1819S, 2012arXiv1207.5303C, 2012Entrp..14.1296G, 2016arXiv160108175D, 1986Ap&SS.123..405B, 1982CaJPh..60..659M, 2007CQGra..24.6243P, 2009PhRvD..79l3509K, 2009PhRvD..80b3501A, 2010PhLB..688..129V, 2012PhRvD..85b3512L, 2012PhRvD..85f4001P, 2014CEJPh..12..123S, 2014PhRvD..90l3538P, 2016IJTP...55...71S, 2007PhLB..656...96B,2004CQGra..21..223E}.  Further, the classic texts \cite{ellis3,waldbook} describe the singularity theorems in significant detail. 

In this paper, we consider a $k=0$ FLRW cosmology with barotropic matter, where $ p = w \mu$, and a positive cosmological constant, $\Lambda > 0$. We demonstrate that assuming only the weak energy condition (WEC), that all such models are past asymptotic to a big bang singularity, thereby extending the classic singularity theorem as described above. Note that, throughout, we use units where $8 \pi G = c = 1$.

\section{The Dynamical Equations}
Following \cite{hervik, ellismac}, we consider an orthonormal frame approach, where one can reduce the Einstein field equations to  a system of first-order, autonomous dynamical equations. Namely, one obtains the Raychaudhuri equation,
\begin{equation}
\label{eq:raych1}
\dot{\theta} + \frac{1}{3}\theta^2 + \frac{1}{2} \left(\mu + 3p \right) - \Lambda = 0,
\end{equation}
the Friedmann equation,
\begin{equation}
\label{eq:fried1}
\frac{1}{3}\theta^2 = \mu + \Lambda,
\end{equation}
and the energy-momentum conservation equation,
\begin{equation}
\label{eq:enmom1}
\dot{\mu} + \theta(\mu + p) = 0.
\end{equation}

Now, the advantage to considering $k=0$ FLRW models, is that one can use the Friedmann equation \eqref{eq:fried1} to eliminate $\mu$ from the Raychaudhuri equation \eqref{eq:raych1}, to obtain a single dynamical equation (after employing a barotropic equation of state, $p = w \mu$):
\begin{equation}
\label{eq:raych2}
\dot{\theta} = -\frac{1}{2} \left(w + 1\right) \left(\theta^2 - 3 \Lambda\right).
\end{equation}

The big bang singularity occurs for when $\theta = \infty$. It is of interest to determine whether $\theta \to \infty$ is indeed a past asymptotic state of such a model. We wish to compactify this infinity by employing the following coordinate transformation, which takes place in the phase space of our ordinary differential equation. Namely, we let
\begin{equation}
\label{eq:transform1}
X = \frac{1}{1 + \exp(-\theta)}.
\end{equation}
Therefore, we have that
\begin{equation}
\lim_{\theta \to \infty} X  = 1.
\end{equation}
Therefore, the singularity $\theta = \infty$ appears as $X = 1$ in our transformed variable.

Further, under the transformation Eq. \eqref{eq:transform1}, the Raychauhduri equation \eqref{eq:raych2} takes the form
\begin{equation}
\label{eq:raych3}
\dot{X} = -\frac{1}{2}(1+w)(X-1)X \left[3\Lambda - \log^2 \left(-1 + \frac{1}{X}\right)\right].
\end{equation}

\section{The Singularity as a Source of Raychaudhuri's Equation}
With the Raychaudhuri equation \eqref{eq:raych3} in hand, we are in a position to describe its fixed points and behaviour with respect to these points. In actuality, we are only interested in the point $X=1$ which corresponds to $\theta = \infty$ in the original coordinates. To prove the existence of a big bang singularity, our goal is to show that the point $X=1$ is a past asymptotic state of Eq. \eqref{eq:raych3}.

The first order of business is to determine whether $X=1$ is indeed a fixed point of Eq. \eqref{eq:raych3}. Indeed, there may be some objection to this because the $\log$ function on the right hand side of Eq. \eqref{eq:raych3} blows up for $X = 1$. However, note that
\begin{equation}
\lim_{X \to 1}  -\frac{1}{2}(1+w)(X-1)X \left[3\Lambda - \log^2 \left(-1 + \frac{1}{X}\right)\right] = 0.
\end{equation}
Therefore, $X=1$ is indeed a fixed point of Eq. \eqref{eq:raych3}. We have compactified the big bang singularity to being a fixed point of our transformed equation. 

What remains to be done is to prove that $X=1$ is a past asymptotic state of the dynamics represented by Eq. \eqref{eq:raych3}. The standard methodology is to first determine local behaviour, that is, whether $X=1$ is a source of the system by analyzing the the sign of the first derivative of the right-hand-side of Eq. \eqref{eq:raych3} at $X=1$. However, this cannot be done in this case. If we denote the right-hand-side of Eq. \eqref{eq:raych3} by $f(X)$, then $f'(X)$ is given by
\begin{dmath}
f'(X) = -\frac{1}{2} (w+1) \left[3 \Lambda  (2 X-1)+(1-2 X) \log ^2\left(\frac{1}{X}-1\right)-2 \log \left(\frac{1}{X}-1\right)\right],
\end{dmath}
which is not defined at $X=1$. Further, 
\begin{equation}
\lim_{X \to 1} f'(X) = -\infty.
\end{equation}
Therefore, a local stability analysis as suggested in \cite{ellis}, for example, will not work in this case.

We therefore make use of Chetaev's instability theorem to show that $X=1$ is unstable, that is, it is a past state of Eq. \eqref{eq:raych3}. Following \cite{arnolddyn}, we note that a differentiable function $f$ is called a \emph{Chetaev function} for a singular point $x_{0}$ of a vector field $v$ if it satisfies the following conditions: the function $f$ is defined on a domain $W$ whose boundary contains $x_{0}$, then $f(x) \to 0$ as $x \to x_{0}, \quad x \in W$; $f > 0$ and $\dot{f} > 0$ everywhere in $W$. Then \emph{Chetaev's instability theorem} states that a singular point for which a Chetaev function exists is unstable.

Motivated by this theorem, let us define a function $Z$, such that
\begin{equation}
Z = -\log X,
\end{equation}
and a domain $W$ such that
\begin{equation}
W = \left\{\frac{1}{1 + \exp \left(-\sqrt{3} \sqrt{\Lambda}\right)} < X  < 1\right\}.
\end{equation}
Clearly, the boundary of $W$ is given by
\begin{equation}
\partial W = \left\{X = \frac{1}{1 + \exp \left(-\sqrt{3} \sqrt{\Lambda}\right)}\right\} \cup \left\{X = 1 \right\},
\end{equation}
which importantly contains the fixed point $X = 1$. 

To apply Chetaev's theorem, first note that
\begin{equation}
\lim_{X \to 1} Z = \lim_{X \to 1} -\log X = 0.
\end{equation}
Further, the function $Z$, and its derivative, $\dot{Z} = -\dot{X}/X$ are clearly strictly positive in $W$ (which can be seen using Eq. \eqref{eq:raych3}) on the condition that $-1 < w \leq 1$ and $\Lambda > 0$. For a barotropic equation of state the condition $w > -1$ is just the weak energy condition. The condition $\Lambda > 0$ indicates that we must have a positive cosmological constant. 

Therefore, we have just shown that for a positive cosmological constant and assuming that the matter distribution in our model satisfies the weak energy condition, $X=1$ is indeed a source of Eq. \eqref{eq:raych3}. That is, there exists a one-dimensional unstable manifold of $X=1$ that is tangent to the one-dimensional unstable subspace at $X=1$ such that all orbits in the unstable manifold are asymptotic to $X=1$ as $t \to -\infty$. 

What remains to be shown is whether $X=1$ is indeed asymptotically unstable. To accomplish this, we will use the LaSalle invariance principle \cite{ellis}, but modified for $\alpha$-limit sets. This reads as follows. Consider a dynamical system $x' = f(x)$ on $\mathbb{R}^{n}$, with flow  $\phi_{t}$. Let $S$ be a closed, bounded, and negatively invariant set of $\phi_{t}$, and let $Z$ be a $C^{1}$ monotone function, where $\dot{Z} \leq 0$. Then, for all $x_{0}$ in $S$, $\alpha(x_{0}) \subseteq \{x \in S | \dot{Z} = 0\}$.  Let $Z_{1} = X$. Then, $Z_1$ is clearly monotone for $X=1$ which is a local source by the preceding analysis, and is therefore a negatively invariant closed set. Therefore, applying the LaSalle invariance principle, we find that $\alpha(x) = \{X=1\}$. We therefore conclude that $X=1$ is asymptotically unstable.

In other words, since $X=1$ represents $\theta = \infty$ in the original variable choice, we have just shown that \emph{a $k=0$ FLRW model with a positive cosmological constant and barotropic matter obeying only the weak energy condition is past asymptotic to a big bang singularity.}

\section{Conclusions}
In this brief paper, we  considered the dynamics of a spatially flat FLRW spacetime with a positive cosmological constant and matter obeying a barotropic equation of state. By performing a coordinate transformation in the configuration space on the Raychaudhuri equation in order to compactify the big bang singularity to a finite point. The fundamental singularity theorem of FLRW cosmologies assumes that the matter content in the cosmological model obeys the strong energy condition along with a nonpositive cosmological constant which gives rise to an irrotational geodesic singularity. We showed that a spatially flat Friedmann-Lema\^{i}tre-Robertson-Walker universe with barotropic matter obeying only the \emph{weak} energy condition with a \emph{nonnegative} cosmological constant also contains a past singularity. This was demonstrated by first showing that the big bang singularity point was locally unstable using Chetaev's instability theorem. We then used the LaSalle invariance principle to show that this point is in fact an asymptotically unstable equilibrium point of Raychaudhuri's equation as applied to the model under consideration.

\section{Acknowledgements}
The author would like to thank Michael C. Haslam, Frits Veerman and Juli\'{a}n Aguirre for helpful discussions regarding fixed points of dynamical systems.

\bibliographystyle{ieeetr} 
\bibliography{sources}

\end{document}